OpenStudy (anonymous):

Find the first 6 terms of each sequence: I will write down in equation editor.

OpenStudy (anonymous):

first time I've seen you here

OpenStudy (anonymous):

\[a _{1}=1, a _{2}=2, a _{n+2}=4a _{n+1}-3a _{n}\]

OpenStudy (anonymous):

Sorry, i take forever on this equation thingie.

OpenStudy (anonymous):

ok, let's find \[a_3\] together

OpenStudy (anonymous):

Okie dokie!

OpenStudy (karatechopper):

its like lil emu is in dsiguise helpin ya!;D lol i will leave ya alone now to get bak to hw

OpenStudy (anonymous):

Okay. Got it.

OpenStudy (anonymous):

\[a_3=a_{1+2}\] which mean n=1 using equation \[a_{2+1}=4a_{1+1}-3a_{1}\]

OpenStudy (anonymous):

wait..so n does not equal 2?

OpenStudy (anonymous):

no, I meant n=1

OpenStudy (anonymous):

Okie.

OpenStudy (anonymous):

\[a_{2+1}=4a_{1+1}−3a_1\] \[=a_{3}=4a_{2}−3a_1\] you know what a1 and a2 are

OpenStudy (anonymous):

Oh coolio!!! We can plug those numbers in.

OpenStudy (anonymous):

a_3=4(2)-3(1)=5

OpenStudy (anonymous):

let see you find a_4

OpenStudy (anonymous):

Okay, i got 5 for a_3

OpenStudy (anonymous):

Would i have to make n equal to 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got 14 for a_4

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

YAAAYYY!!! now its soo simple for me!!

OpenStudy (anonymous):

I got this, thank you thank you!!!!

OpenStudy (karatechopper):

done yet ??????????????